OpenStudy (anonymous):

Powers and products - Help 1. Y4 2. 3X3 - i mean 3 Y to the power 3 3. 2Y2 4. Z2 5. ( YZ)2 6. ( YX) 2 7. X2Z2 8. Yx 9. -Yx Plzz show your work. thanks guys plzz help

OpenStudy (anonymous):

Plzzz someone help!!!!!!!!!!!!!!!!!!!!

OpenStudy (ranga):

What is the question?

OpenStudy (anonymous):

1. Y4 2. 3X3 - i mean 3 Y to the power 3 3. 2Y2 4. Z2 5. ( YZ)2 6. ( YX) 2 7. X2Z2 8. Yx 9. -Yx

OpenStudy (ranga):

You have given a list of things but you are not telling what needs to be done. There are no questions to answer.

OpenStudy (anonymous):

i need help solve them. like #1 Whats Y to the 3 power??

OpenStudy (anonymous):

I mean 4

OpenStudy (ranga):

Solving requires equations. None of the things listed are equations. But if you are asking for an explanation of what this means:\[\Large Y^4\]then I can help. But without an equation there is nothing to solve here.

OpenStudy (anonymous):

Like okay i already did one. ( 4y )2 = 4 to 2 power x Y2.

OpenStudy (ranga):

At least there was something to do there. But Y^4 means Y * Y * Y * Y Is that what you want?

OpenStudy (anonymous):

yes and what it =

OpenStudy (ranga):

If they give you Y = 2 and ask what is Y^4 then we can find that. But asking what is Y^4 does not make much sense. All I can say is Y^4 is Y multiplying itself 4 times

OpenStudy (anonymous):

So just put 4

OpenStudy (ranga):

Maybe you did not copy the full question or maybe the question is a followup on a previous question or something, But the question is incomplete and I am sorry there is nothing much I can do here.

OpenStudy (anonymous):

All the question are evaluate each monomial for x = 5, Y= = -1, and Z = -4

OpenStudy (ranga):

Oh my God! You leave out the essential part of the question and I have wasted SO much time trying to figure out what it is you want!!!

OpenStudy (anonymous):

Sorry lol

OpenStudy (ranga):

Okay, I will do a couple and you do the rest. x = 5, Y= = -1, and Z = -4 1) Y^4 Put Y = -1 Y^4 = (-1)^4 = (-1) * (-1) * (-1) * (-1) = 1 5) (YZ)^2 put Y = -1 and Z = -4 ( -1 * -4 )^2 = (4)^2 = 4 * 4 = 16 7) X^2Z^2 put X = 5, Z = -4 (5)^2 * (-4)^2 = (5 * 5) * (-4 * -4) = 25 * 16 = 400 You do the rest in the same manner.

OpenStudy (anonymous):

Can you plzz do the rest its gonna take me forever

OpenStudy (ranga):

The idea is for you to learn. I think I have given enough examples and you should learn to do it. Initially it may take long as is always the case when learning but eventually it will get easier.

OpenStudy (anonymous):

At less do 2 more plzz.